OpenStudy (anonymous):

Mary's paycheck was for $600.00. When she cashed her check at the bank, the teller gave her a total of 40 bills, both tens and twenties. How many tens did the teller give Mary?

OpenStudy (texaschic101):

x = tens y = twenties x + y = 40 --> y = 40 - x 10x + 20y = 600 now sub 40 - x in for y in the 2nd equation and solve for x.